# jmstate, a Flexible Python Package for Multi-State Joint Modeling

### v1

@@ Abstract
Classical joint modeling approaches often rely on competing risks or recurrent event formulations to account for complex real-world processes involving evolving longitudinal markers and discrete event occurrences. However, these frameworks typically capture only limited aspects of the underlying event dynamics.
Multi-state joint models offer a more flexible alternative by representing full event histories through a network of possible transitions, including recurrent cycles and terminal absorptions, all potentially influenced by longitudinal covariates.
In this paper, we propose a general framework that unifies longitudinal biomarker modeling with multi-state event processes defined on arbitrary directed graphs. Our approach accommodates both Markovian and semi-Markovian transition structures, and extends classical joint models by coupling nonlinear mixed-effects longitudinal submodels with multi-state survival processes via shared latent structures.
We derive the full likelihood and develop scalable inference procedures based on stochastic gradient descent. Furthermore, we introduce a dynamic prediction framework, enabling individualized risk assessments along complex state-transition trajectories.
To facilitate reproducibility and dissemination, we provide an open-source Python library jmstate implementing the proposed methodology, available on PyPI. Simulation experiments and a biomedical case study demonstrate the flexibility and performance of the framework in representing complex longitudinal and multi-state event dynamics. The full Python notebooks used to reproduce the experiments as well as the source code of this paper are available on GitLab.
Keywords joint modeling • multi-state processes • longitudinal data • survival analysis • stochastic gradient descent • dynamic prediction • Python library
@@ 1 Introduction
Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [Papageorgiou et al., 2019], particularly for dynamic prediction in clinical applications. Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [Rizopoulos, 2012], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures. However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [Król et al., 2017]. In such settings, multi-state models provide a more flexible approach [Ferrer et al., 2016].
Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [Lovblom et al., 2024]. In biostatistics, these models are widely used for survival and reliability analysis, allowing for a richer and more accurate representation by capturing alternative paths to an event of interest, intermediate events, and progressive disease. Key components of multi-state models include transition intensity functions, which denote the instantaneous risk of moving from one state to another, and transition probability functions, which describe the probability of transition over longer intervals. Often, these models assume the Markov property, where future transitions depend only on the current state, simplifying the transition intensity functions [Asanjarani et al., 2022].
However, semi-Markov processes extend this by allowing future probability transitions to depend on the sojourn time, with the clock resetting after each transition.
@@
The model proposed by Ferrer et al. [Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016] exemplifies this link by presenting a joint model for a longitudinal process (e.g., Prostate-Specific Antigen measurements) and a multi-state process (e.g., clinical progressions in prostate cancer). These two sub-models are interconnected by shared random effects, allowing the model to account for the correlation between the continuous longitudinal biomarker trajectory and the discrete transitions between health states.
Several software packages have been developed in R and Python to fit joint models that combine longitudinal and time-to-event data. The JMbayes package [Rizopoulos, 2016] offers a Bayesian implementation of joint models using Markov chain Monte Carlo (MCMC) methods. Its successor, JMbayes2 [Rizopoulos et al., 2024], extends this framework to accommodate more complex data structures, such as multiple longitudinal outcomes, nonlinear trajectories, and competing risks. Another popular R package, joineR [Philipson et al., 2018], enables joint modeling of repeated measurements and survival outcomes, based on the methodology introduced by Williamson et al. [2008]. In the Python ecosystem, the flash package [Nguyen et al., 2024] provides a scalable joint modeling approach tailored to high-dimensional settings, exploiting sparse association patterns between longitudinal and survival processes.
Despite the richness of these tools, none currently supports joint models in the multi-state framework with nonlinear longitudinal modeling. In this work, we build on the approach of Ferrer et al. [Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016] and introduce jmstate, a Python package for joint modeling of nonlinear longitudinal and multi-state time-to-event data. The underlying model of jmstate incorporates a linear mixed-effects submodel for the longitudinal biomarker and a multi-state process for the event history, linked through shared random effects as in [Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016].
The model we propose operates on an arbitrary directed graph, enabling representation of complex state transitions and recurrent events. We derive the full likelihood for the nonlinear joint model, introduce an efficient stochastic approximation inference method, and develop dynamic prediction tools. To illustrate its practical relevance, we conduct both a simulation study and a synthetic biomedical case study using the jmstate PyPI package.
This paper is organized as follows. Section 2 presents background on joint modeling frameworks and multi-state processes. Section 3 introduces the proposed multi-state joint model, detailing its likelihood formulation and inference strategies. Section 4 develops dynamic prediction tools. Section 5 details efficient inference methods based on stochastic approximation. Section 6 provides implementation details. Section 7 exemplifies the use of jmstate though a simulation study while Section 8 illustrates the use of jmstate on data from the PAQUID cohort. Finally, Section 9 concludes with a discussion of the main strengths, limitations, and future directions of this work.
@@ 2 Background
2.1 Notations
Each individual $i$, where $i=1,\ldots,n$, is observed at times $t_{ij}$, where $j=1,\ldots,n_{i}$. $n_{i}$ is the number of observations for an individual $i$. The biomarker values at these times are denoted by $Y_{ij}\in\mathbb{R}^{d}$, where $d$ is the number of biomarkers. Each individual also has a set of covariates $X_{i}\in\mathbb{R}^{k}$, which may include demographic or clinical characteristics.
2.2 Standard Joint Modeling Framework
Joint modeling frameworks are essential to link longitudinal and time-to-event data for more accurate dynamic predictions. The standard framework [Rizopoulos, 2012, 2011] consists of:
  • Mixed-Effects for Longitudinal Data. The longitudinal process $Y_{ij}$ for subject $i$ at time $t_{ij}$ is typically modeled as:
    $Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\psi_{i}=f(\gamma,X_{i},b_{i}),$
    where $\gamma$ are the population parameters, $b_{i}$ are individual random effects, and $\epsilon_{ij}$ is a random measurement error term. It is usually assumed to be Gaussian with zero mean. The functions $h$ and $f$ can be nonlinear, capturing complex within-subject trajectories and between-subject variability. In practice, $h$ is often a linear predictor of the biomarker values, while $f$ may represent a nonlinear function of the random effects.
    Both $h$ and $f$ may be nonlinear, allowing flexible modeling of complex biomarker dynamics. Typical choices include:
    • $h(t,\psi_{i})=\psi_{i1}+\psi_{i2}t+\psi_{i3}t^{2}$ (polynomial growth),
    • $h(t,\psi_{i})=\psi_{i1}\exp(-\psi_{i2}t)$ (exponential decay),
    • $f(\gamma,X_{i},b_{i})=\gamma X_{i}+b_{i}$ [Davidian and Giltinan, 2003], $\gamma+b_{i}$, or $\exp(\gamma+b_{i})$ for log-normally distributed parameters.
  • Cox-Based Survival Modeling. The survival process is described by a hazard function incorporating the current (and possibly historical) longitudinal process and baseline covariates:
    $\lambda_{i}(t\mid X_{i},\mathcal{H}_{i}(t))=\lambda_{0}(t)\exp\left(\alpha\,g(t,\psi_{i})+\beta X_{i}\right),$
    where $\mathcal{H}_{i}(t)$ is the information that would be provided by knowing the "true trajectory" of the longitudinal biomarker process, that is, $\mathcal{H}_{i}(t)=\{(u,h(u,\psi_{i})):0\leq u\leq t\}$ and $g(\cdot)$ links the biomarker trajectory to the hazard, with $\alpha$ et $\beta$ as association parameters. The baseline hazard $\lambda_{0}(t)$ may be parametric or semiparametric with a spline basis representation [Andrinopoulou et al., 2018].
  • Prior distribution. A prior has to be specified for the latent variables $b_{i}$. Here, we assume it to be Gaussian with zero mean: $\mathcal{N}(0,Q)$ but other prior distributions can be considered.
In joint modeling frameworks, the longitudinal process is linked to the time-to-event component through a function $g$ of the underlying trajectory. Several choices for $g$ have been proposed, depending on the hypothesized biological mechanism. Current examples include [Papageorgiou et al., 2019, Mauff, 2023]:
  • $g(t,\psi_{i})=h(t,\psi_{i})$: direct effect of the current biomarker level,
  • $g(t,\psi_{i})=\frac{\partial}{\partial t}h(t,\psi_{i})$: effect of the biomarker slope (rate of change),
  • $g(t,\psi_{i})=\int_{0}^{t}h(w,\psi_{i})\,dw$: cumulative exposure.
This joint framework not only captures the dynamics of longitudinal biomarkers but also leverages their predictive power to inform survival outcomes, enabling dynamic, individualized risk predictions [Rizopoulos, 2011, Ibrahim et al., 2010].
Moreover, in order to derive an expression for the marginal (joint) likelihood, joint modeling typically requires the following independence assumptions [Rizopoulos, 2012]:
Assumption 1.
The latent variables $(b_{i})_{i}$ are independent across individuals.
Assumption 2.
The observed markers $(Y_{ij})_{ij}$ are independent conditionally on $(b_{i})_{i}$.
Assumption 3.
The occurrence times $(T_{i}^{*})_{i}$ are independent conditionally on $(b_{i})_{i}$.
Assumption 4.
The occurrence times $(T_{i}^{*})_{i}$ and the observed markers $(Y_{ij})_{ij}$ are mutually independent conditionally on $(b_{i})_{i}$.
Assumption 5.
The censoring times $(C_{i})_{i}$ are independent and noninformative: conditional on $(b_{i})_{i}$ (and possibly covariates), $(C_{i})_{i}$ is independent of both the event times $(T_{i}^{*})_{i}$ and the longitudinal measurements $(Y_{ij})_{ij}$.
These assumptions facilitate the derivation of the marginal likelihood by allowing the decomposition of the joint distribution into manageable components. They ensure that the latent variables capture all relevant dependencies between the longitudinal and survival processes, enabling efficient estimation and inference.
@@ 2.3 Multi-State Markov Processes
A Markov process (or continuous-time Markov chain) is a specific class of multi-state models in which future transitions between states depend only upon the current state. This Markov property means the process is memoryless. A key consequence is that the duration spent in any state follows an exponential distribution, implying a constant hazard rate for leaving that state [Jackson, 2011, Putter et al., 2007].
However, the Markov assumption can be unrealistic in many real-world applications. For example, in the study of human sleep stages, sojourn times often do not follow an exponential distribution [Dong et al., 2008, Roever et al., 2010], and in the case of chronic diseases such as AIDS, the risk of disease progression can depend on the time elapsed since infection [Andersen and Keiding, 1999]. To address these limitations, semi-Markov processes (SMPs) were introduced, which allow for arbitrary distributions of sojourn times while retaining the Markov property for the embedded discrete-time chain. This flexibility makes SMPs suitable for modeling complex disease progression and patient recovery scenarios [Putter et al., 2007].
@@ 2.4 Multi-State Semi-Markov Processes
In an MSMP, the process is defined by a directed graph $G=(V,E)$, where $V$ represents the set of states and $E$ the set of possible transitions. The transition intensities $\lambda_{k^{\prime}|k}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$. The sojourn time in state $k$ is not restricted to an exponential distribution, allowing for more realistic modeling of the time spent in each state.
MSMPs have been widely applied in a range of disciplines. In reliability, they are used to modeling degradation and repair processes [Limnios and Oprisan, 2001]. In biomedical studies, they have proven useful for analyzing illness-death models or disease progression with non-exponential transitions [Commenges et al., 2006, Fiocco et al., 2008]. Applications in finance also exist, where semi-Markovian dynamics can model credit rating migrations or economic regimes [Luciano and Vasiliev, 2006]. In these contexts, MSMPs retain the Markov property in the embedded jump chain while offering increased realism through flexible dwell time modeling.
From a methodological standpoint, MSMPs extend estimation strategies developed for Markov models, such as maximum likelihood or Bayesian inference, and can accommodate interval-censored or misclassified data [Frydman, 2005]. This makes them a powerful and general tool for multi-state event history analysis.
@@ 2.5 Joint Models and Multi-State Processes
Previous research has initiated such integration, with models like the one proposed by Ferrer et al. [Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016], which combines a linear mixed sub-model for longitudinal data with a multi-state sub-model using shared random effects. This model enables the simultaneous analysis of repeated measurements of a biomarker (the longitudinal process, e.g., PSA levels) and the times of transitions between multiple health states (the multi-state process). It explicitly accounts for the link between these two correlated processes and uses information from the biomarker dynamics to explain or predict clinical progression events. The multi-state sub-model assumes a non-homogeneous Markov process for the clinical progression, meaning that the future of the process depends only on the current state (Markov property) and the time elapsed since study entry (non-homogeneous property).
@@
Although Ferrer et al. [Ferrer et al., 2016] note that their joint modeling framework could be adapted to semi-Markov processes in other contexts, their primary application to prostate cancer progression assumes a non-homogeneous Markov multi-state process. Consequently, the model does not explicitly incorporate the effect of time spent within a particular state on the instantaneous risk of transitioning out of that state, beyond the overall follow-up time from study entry.
While this prior work has managed to combine multi-state and joint models in particular applications, these models typically rely on a linear mixed model for the biomarker and impose a specific directed acyclic graph (DAG) structure through their likelihood decomposition, which restricts the types of dependency structures they can represent. As a result, limitations remain in their generalizability, flexibility, and the scalability of inference. In what follows, we propose a more general framework defined on arbitrary directed graphs (not restricted to acyclic forms) that accommodates both Markov and semi-Markov assumptions and allows nonlinear longitudinal submodels.
3 Extending Multi-State Joint Models
@@
Joint models traditionally focus on a single time-to-event outcome, which limits their ability to represent complex disease trajectories involving multiple intermediate and recurrent events. To address this limitation, we extend the classical nonlinear joint modeling framework to incorporate multi-state processes, enabling the joint analysis of longitudinal biomarkers and transitions between multiple health states. This is achieved by introducing a directed graph structure $G=(V,E)$, which encodes all permissible transitions and supports both Markov and semi-Markov assumptions.
3.1 Graph Structure and Example
Let us consider the illustrative state-transition diagram shown in Figure 1, representing a 4-state process.
Figure 1: Illustration of the 4-state transition graph $G=(V,E)$, including absorbing state 4 (Dead).
This graph can be encoded by an adjacency matrix $A=[A_{kk^{\prime}}]_{kk^{\prime}}$ where $A_{kk^{\prime}}=1$ if a transition from state $k$ to $k^{\prime}$ is possible, and 0 otherwise. For the graph in Figure 1, the adjacency matrix is:
$A=\begin{pmatrix}0&1&0&1\\ 0&0&1&1\\ 0&1&0&1\\ 0&0&0&0\end{pmatrix}$
Under certain assumptions borrowed from the semi-Markov literature, competing and recurrent risks mentioned in the introduction, as well as the classical joint model, can be represented using such graphs as illustrated in Figure 2.

### v2

# A General Framework for Joint Multi-State Models

@@ Abstract
Classical joint modeling approaches often rely on competing risks or recurrent event formulations to describe complex processes involving evolving longitudinal biomarkers and discrete event occurrences, but these frameworks typically capture only limited aspects of the underlying event dynamics.
We propose a general multi-state joint modeling framework that unifies longitudinal biomarker dynamics with multi-state time-to-event processes defined on arbitrary directed graphs. The proposed framework accommodates arbitrary directed transition graphs, nonlinear longitudinal submodels, and scalable inference via stochastic gradient descent. This formulation encompasses both Markovian and semi-Markovian transition structures, allowing recurrent cycles and terminal absorptions to be naturally represented. The longitudinal and event processes are linked through shared latent structures within nonlinear mixed-effects models, extending classical joint modeling formulations.
We derive the complete likelihood, establish conditions for identifiability, and develop scalable inference procedures based on stochastic gradient descent to enable high-dimensional and large-scale applications. In addition, we formulate a dynamic prediction framework that provides individualized state-transition probabilities and personalized risk assessments along complex event trajectories.
Through simulation and application to the PAQUID cohort, we demonstrate accurate parameter recovery and individualized prediction.
keywords:
joint modeling, multi-state processes, longitudinal data, survival analysis, stochastic gradient descent, dynamic prediction
@@ 1 Introduction
Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [papageorgiou2019overview], particularly for dynamic prediction in clinical applications. Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [rizopoulos2012joint], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures. However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [KrolMauguen2017]. In such settings, multi-state models provide a more flexible approach [ferrer2016joint].
Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [lovblom2024modeling]. In biostatistics, these models are widely used for survival and reliability analysis, allowing for a richer and more accurate representation by capturing alternative paths to an event of interest, intermediate events, and progressive disease. Key components of multi-state models include transition intensity functions, which denote the instantaneous risk of moving from one state to another, and transition probability functions, which describe the probability of transition over longer intervals. Often, these models assume the Markov property, where future transitions depend only on the current state, simplifying the transition intensity functions [asanjarani2022estimation].
The Markov assumption can be relaxed by adopting a semi-Markov formulation, in which the transition probabilities depend not only on the current state but also on the sojourn time, effectively resetting the time scale after each transition.
@@
The model proposed by Ferrer et al. [ferrer2016joint] exemplifies this link by presenting a joint model for a longitudinal process (e.g., Prostate-Specific Antigen (PSA) measurements) and a multi-state process (e.g., clinical progressions in prostate cancer). These two sub-models are interconnected by shared random effects, allowing the model to account for the correlation between the continuous longitudinal biomarker trajectory and the discrete transitions between health states.
The model we propose operates on an arbitrary directed graph, enabling representation of complex state transitions and recurrent events. We derive the complete likelihood for the nonlinear joint model, introduce an efficient stochastic approximation inference method, and develop dynamic prediction tools. To illustrate its practical relevance, we conduct both a simulation study and a synthetic biomedical case study.
This paper is organized as follows. Section 2 reviews the background on multi-state and joint modeling frameworks. Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and identifiability conditions. Section 4 describes the stochastic-gradient–based inference algorithm, while Section 5 presents the dynamic prediction framework. Section 6 reports simulation experiments that assess parameter recovery, identifiability, and convergence properties of the proposed model, and Section 7 applies the framework to data from the PAQUID cohort. Finally, Section 8 concludes with perspectives and future research directions.
@@ 2 Background
Joint modeling of longitudinal and time-to-event data provides a unified statistical framework to analyze the interplay between continuous biomarker dynamics and event occurrence processes. In its classical form, this framework couples a mixed-effects model describing individual biomarker trajectories with a survival model for event times [rizopoulos2012joint]. Such models have become central in biomedical research, particularly for dynamic prediction and personalized risk assessment [papageorgiou2019overview]. However, these traditional joint models generally focus on a single terminal event, limiting their ability to represent more complex event histories that include intermediate, recurrent, or competing events. To address these limitations, multi-state extensions have been developed, providing a natural framework for describing transitions between multiple clinical or biological states over time [ferrer2016joint, KrolMauguen2017].
@@ 2.1 Multi-State Markov Processes
A Markov process (or continuous-time Markov chain) is a specific class of multi-state models in which future transitions between states depend only upon the current state. This Markov property means the process is memoryless. A key consequence is that the duration spent in any state follows an exponential distribution, implying a constant hazard rate for leaving that state [jackson2011multi, putter2007competing].
However, the Markov assumption can be unrealistic in many real-world applications. For example, in the study of human sleep stages, sojourn times often do not follow an exponential distribution [dong2008non, roever2010statistical], and in the case of chronic diseases such as AIDS, the risk of disease progression can depend on the time elapsed since infection [andersen1999multi]. To address these limitations, semi-Markov processes (SMPs) were introduced, which allow for arbitrary distributions of sojourn times while retaining the Markov property for the embedded discrete-time chain. This flexibility makes SMPs suitable for modeling complex disease progression and patient recovery scenarios [putter2007competing].
@@ 2.2 Multi-State Semi-Markov Processes
In an MSMP, the process is defined by a directed graph $G=(V,E)$, where $V$ represents the set of states and $E$ the set of possible transitions. The transition intensities $\lambda_{k^{\prime}\mid k}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$ are therefore assumed to be non-identically zero. The sojourn time in state $k$ is not restricted to an exponential distribution, allowing for more realistic modeling of the time spent in each state.
MSMPs have been widely applied in a range of disciplines. In reliability, they are used to modeling degradation and repair processes [limnios2001semi]. In biomedical studies, they have proven useful for analyzing illness-death models or disease progression with non-exponential transitions [commenges2006likelihood, fiocco2008modelling]. Applications in finance also exist, where semi-Markovian dynamics can model credit rating migrations or economic regimes [luciano2006semi]. In these contexts, MSMPs retain the Markov property in the embedded jump chain while offering increased realism through flexible dwell time modeling.
From a methodological standpoint, MSMPs extend estimation strategies developed for Markov models, such as maximum likelihood or Bayesian inference, and can accommodate interval-censored or misclassified data [frydman2005estimation]. This makes them a powerful and general tool for multi-state event history analysis.
@@ 2.3 Joint Models and Multi-State Processes
Previous research has initiated such integration, with models like the one proposed by Ferrer et al. [ferrer2016joint], which combines a linear mixed sub-model for longitudinal data with a multi-state sub-model using shared random effects. This model enables the simultaneous analysis of repeated measurements of a biomarker (the longitudinal process, e.g., PSA levels) and the times of transitions between multiple health states (the multi-state process). It explicitly accounts for the link between these two correlated processes and uses information from the biomarker dynamics to explain or predict clinical progression events. The multi-state sub-model assumes a non-homogeneous Markov process for the clinical progression, meaning that the future of the process depends only on the current state (Markov property) and the time elapsed since study entry (non-homogeneous property).
@@
Although Ferrer et al. [ferrer2016joint] note that their joint modeling framework could be adapted to semi-Markov processes in other contexts, their primary application to prostate cancer progression assumes a non-homogeneous Markov multi-state process. However, this framework is not well suited for arbitrary transition patterns other than the ones modeled by DAGs.
While both joint models and multi-state models offer powerful tools for analyzing disease progression, they emphasize different aspects of the underlying process. Multi-state models capture the sequence and timing of discrete transitions between states, whereas joint models focus on the evolution of continuous longitudinal markers and their association with event risk. Existing joint multi-state models, such as that of Ferrer et al. [ferrer2016joint], are restricted to linear mixed-effects submodels and to acyclic Markov transition graphs, as their likelihood relies on a sequential factorization that is not valid when transitions form cycles. However, many biological processes involve cyclic or semi-Markov transitions and nonlinear biomarker dynamics. These cannot be handled by existing frameworks, motivating the present general formulation.
In contrast, many real-world processes—such as recurrent disease relapse or recovery—require more flexible representations involving nonlinear biomarker dynamics, semi-Markovian dependencies, and recurrent cycles. The next section introduces a general framework that integrates these features within a unified likelihood-based formulation, enabling efficient inference and dynamic prediction along complex event trajectories.
3 A General Multi-State Joint Modeling Framework
@@
Classical joint models typically link a longitudinal biomarker process to a single time-to-event outcome. While effective for terminal events, this design cannot represent more complex disease trajectories involving intermediate or recurrent events. To address these limitations, we extend the nonlinear joint modeling framework to incorporate a multi-state process, thereby enabling a unified analysis of longitudinal biomarkers and discrete transitions between multiple health states.
The proposed model defines a joint process $\{Y_{i}(t),S_{i}(t)\}_{t\geq 0}$, where $Y_{i}(t)$ denotes the longitudinal biomarker trajectory and $S_{i}(t)$ the multi-state event process. Both components are linked through shared latent structures, and the event process is represented on an arbitrary directed graph supporting both Markovian and semi-Markovian transition mechanisms.
3.1 Notation Conventions
Throughout this section, $i=1,\dots,n$ indexes individuals, and $t_{ij}$ denotes the $j$-th observation time of individual $i$. The vector $Y_{ij}\in\mathbb{R}^{d}$ represents longitudinal biomarker measurements, and $X_{i}\in\mathbb{R}^{p}$ the associated covariates.
The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, while $\psi_{i}=f(\gamma,X_{i},b_{i})$ denotes the subject-specific parameters in the longitudinal submodel. The filtration $\mathcal{H}_{i}(t)$ represents the observed history of the longitudinal process up to time $t$.
Transition times and states are denoted by $(T_{i\ell},S_{i\ell})$, with right-censoring time $C_{i}$. We also define the index of the last observed transition before time $t$ denoted by $m_{i}(t)\coloneq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$.
3.2 Extending Multi-State Joint Models
Joint models traditionally focus on a single time-to-event outcome, which limits their ability to represent complex disease trajectories involving multiple intermediate and recurrent events. To address this limitation, we extend the classical nonlinear joint modeling framework to incorporate multi-state processes, enabling the joint analysis of longitudinal biomarkers and transitions between multiple health states. This is achieved by introducing a directed graph structure encoding all permissible transitions and supports both Markov and semi-Markov assumptions.
3.2.1 Graph Structure
Let $G=(V,E)$ be a directed graph, where $V=\{1,\dots,p\}$ denotes the set of states and $E\subset V\times V$ the set of admissible transitions. The graph encodes all possible paths of the multi-state process, allowing for competing, recurrent, or absorbing transitions. Figure 1 illustrates an example of a four-state model including an absorbing state (Death). The corresponding adjacency matrix $A=(A_{kk^{\prime}})_{1\leq k,k^{\prime}\leq|V|}$ satisfies $A_{kk^{\prime}}=1$ if $(k,k^{\prime})\in E$ and 0 otherwise.
Figure 1: Example of a 4-state transition graph $G=(V,E)$, including an absorbing state (4).
Such a representation generalizes standard joint models, which can be recovered as special cases: single-event (linear chain), competing-risks (two absorbing transitions), or recurrent-event settings (cyclic transitions) as illustrated in Figure 2.